# Mask of truth: model sensitivity to unexpected regions of medical images

### v2

@@ 1 Introduction
    We compare the embeddings and apply explainability methods to better understand which elements are used to perform the classification.
@@ 2 Related Work
Recent studies show that datasets can suffer from problems like label errors, biases or shortcuts in datasets which may impact the reliability of the evaluation. In this work, we focus on shortcuts, also called spurious correlations or confounders, leading models to be biased and use irrelevant information to make decisions [22, 23].
@@
Some, more similar to our method, study the performance of models when removing the area of interest. Bissoto et al. [28] perform the analysis of melanoma classification removing the lesion in the image. Hemelings et al. [29] also evaluate the performances of models on eye fundus photographies under different masking strategies showing good results despite removing the optic disc. Both studies only evaluate models on the same type of images (e.g. train and evaluate without the ROI) and did not assess whether clinicians were able to diagnose the images correctly. Beyond classification models, shortcuts learning is also a phenomenon impacting segmentation models as highlighted by Lin et al. [30].
@@
Our study adds to this line of research by analysing more recent datasets and including a domain expert to strengthen our findings with clinical knowledge. Moreover, our method doesn’t rely on prior annotation of potential shortcuts but on the knowledge of clinically relevant parts of the image.
@@ 3.4 Explainability method
We use SHAP (SHapley Additive exPlanations) [46] following the results of Sun et al. [34] showing its better performance on images containing shortcuts compared to gradient-based explanation methods such as Grad-Cam [47]. SHAP divides the image into superpixels and computes the fair contribution of each superpixel using the change in the classifier outputs after their alteration using occlusion, blurring or inpainting method. We used an occlusion mask and set the number of evaluations to 1000.
@@ 3.5 Study with a domain expert
We compare the ability of the AI models and a domain expert to distinguish healthy and non-healthy images to assess whether it was possible to make the correct diagnosis for clinically relevant reasons. We asked a radiology resident with five years of experience and who has been involved previously in multiple AI projects to annotate images from each condition and the different masking strategies. The study was performed in two steps. During the first step, the radiology resident had to annotate ten images randomly selected from our training set and could ask any question. This phase was used to present the study to the radiology resident, to ask his opinion on the annotation tool and to improve it.
@@ 4.1 Different impact between chest X-rays and eye fundus
On the PadChest dataset, we found that every model is able to obtain good performances when trained and evaluated on similar masking even on images without lungs using the bounding box which removes a high proportion of the image (AUC between 0.85 and 0.93 on the effusion class and minimum 0.74 on the pneumonia with models trained on No lungs BB).
Another interesting result is that the model trained on Full images performed better when evaluated on images without lungs (0.86 AUC) than on images with only lungs (0.47 AUC) indicating that the model is not able to distinguish images with effusion when keeping only the lungs with the precise mask. However, we see an important increase when using the bounding box (0.91 AUC) which may indicate the usage of elements near the lungs captured when using the bounding box. These results, especially the surprisingly high performances of models trained on images without lungs using the bounding box, reinforced our hypothesis of models affected by spurious correlation in the dataset.
@@
We evaluate the evolution of AUC when increasing the size of the masks for both Only disc and No disc. In this experiment, we either only increase the size of the masks for healthy images or glaucoma images.
@@
Figure 5: Evolution of AUC when increasing mask size of healthy images (in blue) and glaucoma images (in orange)
@@ 4.2 Poor out-of-distribution performance
We hypothesised that the models trained without the region of interest would have lower performances on external datasets as shortcuts are often dataset-specific, relying on such shortcuts would therefore not transfer well to other datasets. Results in Table 2 seem to partly align with our intuition. We indeed find poor performances of models trained without the ROI, almost always close to an AUC of 0.5. However, we can see that the results of models trained with only the ROI are often also close to this value. In most cases, the best performances are achieved by models trained on Full images (cardiomegaly, effusion, pneumonia and glaucoma), but remain well below the results on the training dataset. This result confirms the poor transferability across populations and the need to test and fine-tune models on the target population before applying the models in a different context like a different hospital.
@@
Only ROI 0.54+/-0.013 0.58+/-0.022 0.53+/-0.021 0.55+/-0.014 0.60+/-0.029 0.53+/-0.031
Table 2: Out-of-distribution evaluation results - Mean AUC (+/- standard deviation) of 5-fold models per masking type and class on full images of the ChestX-ray14 dataset’s test set (Atelectasis, Cardiomegaly, Effusion, Pneumonia and Pneumothorax) and AIROGS dataset (Glaucoma)
@@ 4.3 Closer embeddings between full images and without ROI
For eye fundus, the results in Table 4 are aligned with the t-SNE and show the high proximity of images without the ROI and Full images compared to images with only the ROI. It however doesn’t align with the AUC measure in Fig. 4(b) as we would expect AUC of Full images to be closer to No ROI images than Only ROI images. The close embeddings before the classification head may be caused by the similarity of base images as Full and No lungs images share more common pixels than Full and Only lungs. Therefore, it shows that the result from t-SNE visualisations and cosine similarities of features before the classification head must be interpreted carefully as the classification head can have a major impact.
@@ 4.4 Models can use both the relevant structure and the shortcut
Figure 7: SHAP values of the cardiomegaly class with and without a pacemaker, and without the lungs region. Red values indicate parts increasing the probability
@@ 4.5 Lack of ROI and context makes the diagnosis difficult for a domain expert
Out of the 35 cases of the conditions present in images without the ROI (using the mask or bounding box) only 2 were found by the radiology resident, both in images with lungs removed using the mask. One was a patient with cardiomegaly which is diagnosed using the heart still present when only removing the lungs. This result strengthens our hypothesis that the model relies on non-relevant features to domain experts. Results on full images and images containing only the ROI also reveal the difficulty of the task when only relying on the image. A recurring observation across images is the lack of information about the projection as in Fig. 9. Due to the automatic selection of images based on probabilities, we also found cases in Fig. 10 of quality problems both in the original and the mask dataset. This highlights one of the problems of using large-scale public datasets for which the assessment of quality is difficult.
@@
(a) Image with poor quality
(b) Image with a non-accurate mask
Figure 10: Example of images with either a bad quality making the diagnosis difficult for a radiologist or with a non-accurate mask
@@ 5 Discussion
We studied how convolutional neural networks are affected by the masking of different parts of medical images to highlight the effect of shortcut learning. We analyzed performances, comparing the AUC of each masking strategy, but also went beyond by using an explainability method and comparing the embeddings. Our experiments showed the ability of CNN models to rely on non-relevant parts of medical images. Our results highlight the difficulty of detecting such behaviours. It may sometimes be clear from the non-difference in AUC after the removal of ROI, like in chest X-rays. However, it can also be more hidden like in glaucoma detection for which the performance analysis is not enough. With eye-fundus photography, we also showcased a limitation of relying on the explainability method as the model could be focusing on a correct region but not using the correct feature. It also shows the potential difficulty of detecting patient-level shortcuts described in [31] such as anatomical confounders when they are located at the area of interest.
@@
We focused on the discriminative capacity of models using the AUC as the metric. However, other aspects, such as calibration or fairness across different subgroups, need to be taken into account during the evaluation, especially if the aim is to deploy the model in a clinical environment [54]. Finally, our method only provides a way to detect shortcut learning, other techniques for improving data quality and/or mitigating the model bias would therefore need to be applied.
We believe that the analysis of potential biases should be an important step in the evaluation of AI models. In addition to the poor transferability to other domains such as another hospital, these spurious correlations can have a major impact on the diagnosis of underrepresented populations. We also believe that our field would benefit from going beyond widely used performance metrics such as the AUC that don’t necessarily translate into real clinical progress. To this aim, collaboration with hospitals and domain experts is essential to ground research in more clinical knowledge and ensure its relevance and added value from a clinical point of view.
@@ 5.0.1 Acknowledgments
We thank Casper Anton Poulsen, Trine Naja Eriksen and Cathrine Damgaard for their early work within this research line. This work was supported by the DFF - Inge Lehmann 1134-00017B and Novo Nordisk Foundation NNF21OC0068816

### v3

@@ Abstract
This version of the article has been accepted for publication, after peer review but is not the Version of Record and does not reflect post-acceptance improvements, or any corrections. The Version of Record is available online at: http://dx.doi.org/10.1007/s10278-025-01531-5.
@@ 1 Introduction
    We use embedding analysis and an explainability method to better understand which elements are used but also discuss the limitations of such techniques in this setting.
@@ 2 Related Work
Recent studies show that datasets can suffer from problems like label errors, biases or shortcuts in datasets which may impact the reliability of the evaluation. In this work, we focus on shortcuts, also called spurious correlations or confounders, leading models to be biased and use irrelevant information to make decisions [22, 23, 24].
@@
Some, more similar to our method, study the performance of models when removing the area of interest. Bissoto et al. [29] perform the analysis of melanoma classification removing the lesion in the image. Hemelings et al. [30] also evaluate the performances of models on eye fundus photographies under different masking strategies showing good results despite removing the optic disc. Both studies only evaluate models on the same type of images (e.g. train and evaluate without the ROI) and did not assess whether clinicians were able to diagnose the images correctly. Haynes et al. [31] assess the capacity of Covid-19 classifiers to correctly classify images without the lungs. They find that models, especially when not pre-trained on related tasks, are still able to distinguish healthy and non-healthy images. Aslani et al. [32] present a pipeline improving COVID-19 detection in chest X-rays. The pipeline includes the segmentation of the lungs and crops the image to centralize the lungs, removing some possible biases on the image’s borders such as texts. Beyond classification models, shortcut learning is also a phenomenon impacting segmentation models as highlighted by Lin et al. [33].
@@
Our study adds to this line of research by analysing more recent datasets in two applications and including a domain expert to strengthen our findings with clinical knowledge. Moreover, our method doesn’t rely on prior annotation of potential shortcuts but on the knowledge of clinically relevant parts of the image.
@@ 3.2 Experimental setting, masking of ROI and evaluation
As a baseline, we train logistic regression classifiers using tabular data for the Padchest dataset using the same folds as for the image classification models but removing the samples for which at least one of the features was absent. For the chest x-rays, we train the models with the patient’s birth year, the patient’s sex and the projection view. No patient metadata was available for the eye fundus dataset to conduct such experiments. Baseline models on tabular data are recommended to evaluate the additive value of images compared to metadata [48].
We evaluated the statistical significance of the differences between the AUCs of the different models with the Delong test [49, 50]. The Delong test is the most appropriate test to compare correlated AUC and was for example used in [51, 52] to compare classifiers on medical images. As we have five models per version of the dataset, we concluded that a model was significantly better if the p-values were below 0.05 for at least three folds.
@@ 3.4 Explainability method
We use SHAP (SHapley Additive exPlanations) [54] which divides the image into superpixels and computes the fair contribution of each superpixel using the change in the classifier outputs after their alteration using occlusion, blurring or inpainting method. We used an occlusion mask and set the number of evaluations to 1000.While other works pointed some limitations to its usage [55, 56], we decided to use SHAP following the results of Sun et al. [37]. They show its better performance on medical images containing shortcuts compared to gradient-based explanation methods such as Grad-Cam [57].
@@ 3.5 Study with a domain expert
We compare the ability of the AI models and a domain expert to distinguish healthy and non-healthy images to assess whether it was possible to make the correct diagnosis for clinically relevant reasons in the images. It is for example useful for the effusion class which appears in the pleural space and could therefore be visible in images without lungs depending on the masks. We asked a radiology resident with five years of experience and who has been involved previously in multiple AI projects to annotate images from each condition and the different masking strategies. The study was performed in two steps. During the first step, the radiology resident had to annotate ten images randomly selected from our training set and could ask any question. This phase was used to present the study to the radiology resident, to ask his opinion on the annotation tool and to improve it.
@@ 4.1 Different impact between chest X-rays and eye fundus
On the PadChest dataset, we found that every model is able to obtain good performances when trained and evaluated on similar masking even on images without lungs using the bounding box which removes a high proportion of the image (AUC between 0.85 and 0.93 on the effusion class and minimum 0.74 on the pneumonia with models trained on No lungs BB). These values are all above the baseline results obtained with the logistic regression classifiers on tabular data (mean AUC: 0.71 atelectasis, 0.77 cardiomegaly, 0.72 effusion, 0.60 pneumonia, 0.70 pneumothorax), it is however interesting to note that these AUCs are also above the random value of 0.5. Another interesting result is that for all the classes, models trained on Full images performed better when evaluated on images without lungs than on images with only lungs. These results show that the models are less able to distinguish images when keeping only the lungs with the precise masks with even some classes near or below the random value (atelectasis, effusion, and pneumonia). However, we see an important increase when using the bounding box which may indicate the usage of elements near the lungs captured when using the bounding box. It is worth noting that depending on the receptive field of the model, features at the boundaries of the lungs in Only lungs images may be impacted by the masking.
Fig 5 shows the evolution of the AUC when increasing the mask size for images with only the lungs and without lungs applied to the model trained of Full images. Fig. 18 in Appendix 0.B shows examples of images obtained at each dilation factor. For most classes, the highest drop of performance for images without lungs happens at a dilation factor of 100 or higher when most of the image is masked out. It is interesting to note that despite a drop, the AUC for the pneumonia class remains above random until completely masking out the images. For images with only the lungs we can see that a dilation factor of 50 is needed before achieving results close to Full images. As we concluded from the AUCs using the bounding boxes, these results may reveal the usage of elements near the lungs but also sometimes at the border of the image.
The combination of these results, consistent over all classes, reinforced our hypothesis of models affected by spurious correlation in the dataset.
@@
(a) Images without lungs
(b) Images with only lungs
Figure 5: Evolution of the AUC with standard deviation of the models trained on Full images applied to (a) images with only lungs and (b) images without the lungs while dilating the mask.
@@
We evaluate the evolution of AUC when increasing the size of the masks for both Only disc and No disc. In this experiment, we either only increase the size of the masks for healthy images or glaucoma images. Fig. 19 in Appendix 0.B shows example of both type of image at different dilation factor.
@@
Figure 6: Evolution of the AUC with standard deviation when increasing mask size of healthy images (in blue) and glaucoma images (in orange)
@@ 4.2 Poor out-of-distribution performance
We hypothesised that the models trained without the region of interest would have lower performances on external datasets as shortcuts are often dataset-specific, relying on such shortcuts would therefore not transfer well to other datasets. Results in Table 2 seem to partly align with our intuition. We indeed find poor performances of models trained without the ROI, almost always close to an AUC of 0.5. However, we can see that the results of models trained with only the ROI are often also close to this value. For cardiomegaly, effusion, pneumonia and glaucoma, the best performances are achieved by models trained on Full images. Moreover, we observe no statistically significant difference between Full and Only lung on atelectasis for 3/5 folds. However, the results remain well below the results on the training dataset. This result confirms the poor transferability across populations and the need to test and fine-tune models on the target population before applying the models in a different context like a different hospital.
@@
Only ROI 0.54+/-0.013 0.58+/-0.022 0.53+/-0.021 0.55+/-0.014 0.60+/-0.029* 0.53+/-0.031
Table 2: Out-of-distribution evaluation results - Mean AUC (+/- standard deviation) of 5-fold models per masking type and class on full images of the ChestX-ray14 dataset’s test set (Atelectasis, Cardiomegaly, Effusion, Pneumonia and Pneumothorax) and AIROGS dataset (Glaucoma). * indicates statistical significance for three or more folds (p-values < 0.05)
@@ 4.3 Closer embeddings between full images and without ROI
For eye fundus, the results in Table 4 are aligned with the t-SNE and show the high proximity of images without the ROI and Full images compared to images with only the ROI. It however doesn’t align with the AUC measure in Fig. 4(b) as we would expect AUC of Full images to be closer to No ROI images than Only ROI images. The close embeddings before the classification head may be caused by the similarity of base images as Full and No lungs images share more common pixels than Full and Only lungs.
These results show that despite being commonly used, the interpretation of t-SNE visualisations and cosine similarities of features before the classification head is limited, should be considered carefully, and completed with further analysis.
@@ 4.4 Models can use both the relevant structure and the shortcut
Figure 8: SHAP values of the cardiomegaly class with and without a pacemaker, and without the lungs region. Red values indicate parts increasing the probability. The blue arrows in (a) show the pacemaker location.
@@ 4.5 Lack of ROI and context makes the diagnosis difficult for a domain expert
Out of the 35 cases of the conditions present in images without the ROI (using the mask or bounding box) only 2 were found by the radiology resident, both in images with lungs removed using the mask. One was a patient with cardiomegaly which is diagnosed using the heart still present when only removing the lungs. This result strengthens our hypothesis that the model relies on non-relevant features to domain experts. Results on full images and images containing only the ROI also reveal the difficulty of the task when only relying on the image. A recurring observation across images is the lack of information about the projection as in Fig. 10.
Due to the automatic selection of images based on probabilities, we also found cases in Fig. 11 of quality problems both in the original and the mask dataset. This highlights one of the problems of using large-scale public datasets for which the assessment of quality is difficult but essential as shown on other tasks and datasets to obtain better and more robust models [61, 62].
@@
(a) Image with poor quality.
(b) Image with a non-accurate mask.
Figure 11: Example of images with either a bad quality making the diagnosis difficult for a radiologist due to the blur in the whole image or with a non-accurate mask indicated by the red box outside the chest area.
@@ 5 Discussion
We studied how convolutional neural networks are affected by the masking of different parts of medical images to highlight the effect of shortcut learning. We analyzed performances, comparing the AUC of each masking strategy, but also went beyond by using an explainability method and comparing the embeddings. Our experiments showed the ability of CNN models to rely on non-relevant parts of medical images. Our results highlight the difficulty of detecting such behaviours. The AUC obtained on different masking configurations may show strong evidence of shortcut learning, like in chest X-rays. However, it can also be more hidden like in glaucoma detection for which the performance analysis reveals no signs of shortcut learning. With eye-fundus photography, we also showcased a limitation of relying on the explainability method as the model could be focusing on a correct region but not using the correct feature. It also shows the potential difficulty of detecting patient-level shortcuts described in [34] such as anatomical confounders when they are located at the area of interest.
@@
We focused on the discriminative capacity of models using the AUC as the metric. However, other aspects, such as calibration or fairness across different subgroups, need to be taken into account during the evaluation, especially if the aim is to deploy the model in a clinical environment [66]. Finally, our method only provides a way to evaluate the risk of shortcut learning, further analysis may be performed to conclude on its absence or presence and other techniques for improving data quality and/or mitigating the model bias would therefore need to be applied.
We believe that the analysis of potential biases should be an important step in the evaluation of AI models. In addition to the poor transferability to other domains such as another hospital, these spurious correlations can have a major impact on the diagnosis of underrepresented populations. We also believe that our field would benefit from going beyond widely used performance metrics such as the AUC that don’t necessarily translate into real clinical progress. To this aim, collaboration with hospitals and domain experts is essential to ground research in more clinical knowledge, ensuring better data understanding but also the relevancy and added value of developed algorithms from a clinical point of view.
Both our and other recent works show the need to question whether models the community trains actually learn salient characteristics of the diseases. It is likely that published models overfit to spurious features, and should not be applied in practice until the community figures out how to resolve these issues and achieve improved robustness.
@@ 5.0.1 Acknowledgments
We thank Casper Anton Poulsen, Trine Naja Eriksen and Cathrine Damgaard for their early work within this research line. We also thank Evangelia Christodoulou for her advice on the statistical significance test. This work was supported by the DFF - Inge Lehmann 1134-00017B and Novo Nordisk Foundation NNF21OC0068816.
@@ Appendix 0.A AUC per label and masking
Appendix 0.B Example of images with different dilation factors
(a) Images without lungs
(b) Images with only lungs
Figure 18: Example of NoLungs and OnlyLungs images while increasing the mask size. In this example, the mask covers the entire image at a dilation factor of 300 or more, but a higher dilation may be needed depending on the original mask.
(a) Images without the optic disc
(b) Images with only the optic disc
Figure 19: Example of NoDisc and OnlyDisc images while increasing the mask size.